# CogniSNN: A First Exploration to Random Graph Architecture based Spiking Neural Networks with Enhanced Expandability and Neuroplasticity

### v1

@@ 5.5 Neuroplasticity
[…]
@@
[…]
@@ 6 Conclusion
This paper validates the feasibility of RGA-based SNNs and develops a CogniSNN with expandability and neuroplasticity using the proposed ResNode, the OR gate-driven skip connection mechanism and the critical path-based LWF algorithm. The results are surprising: Compared to traditional SNNs, the CogniSNN is competitive in image recognition and continual learning. This research opens numerous exciting avenues, including leveraging the expandability of CogniSNN for continual learning by automatically adjusting its structure, or optimizing the graph structure to enhance performance.
Ethical Statement
There are no ethical issues.
Acknowledgments
Contribution statement

### v2

@@ 5.5 Path-Plasticity
Table 5: Comparison of accuracy ($\%$) between dissimilar tasks.
@@
In Table 4, the performance of CogniSNN with fine-tuned parameters associated with the path in high $C_{B}$ is marginally more effective than fine-tuned based on low $C_{B}$, showing an improvement of $0.4\%$. Similarly, the performance of low $C_{B}$ improves by $0.2\%$ than that of high $C_{B}$ in Table 5. The improvement is minor due to the simplicity of the classification task in MNIST and CIFAR10. However, this supports our hypothesis regarding the correlation between the selection of the critical path and the similarity of the task (Definition 2).
Relative to the benchmark, the level of forgetting is reduced. When compared specifically to the LwF-based method, our algorithm demonstrates a $3.6\%$ and $5.1\%$ reduction in forgetting on the previous task. This confirms that the critical path-based LwF can effectively mitigate catastrophic forgetting.
5.6 Energy Analysis
The energy consumption advantage of SNNs over ANNs for the same tasks has been demonstrated in many studies [2025_sun_ilif]. In this paper, we propose a spiking residual operation of the OR Gate to reduce the need for real-valued calculations, as well as a complex SNN architecture where ResNodes are randomly connected. Therefore, it is essential to analyze their energy consumption.
Table 6: The total energy consumption for different tasks.
T CogniSNN CogniSNN chain-like SNN
(OR) (ADD) (OR)
DVS-Gesture 5 7.46 mJ 8.89 mJ 7.75 mJ
8 7.86 mJ 9.02 mJ 7.91 mJ
16 8.61 mJ 9.72 mJ 8.43 mJ
CIFAR10-DVS 5 133.62 mJ 135.32 mJ 132.11 mJ
8 118.92 mJ 126.07 mJ 115.13 mJ
N-Caltech101 5 113.69 mJ 115.94 mJ 110.88 mJ
8 101.53 mJ 114.15 mJ 98.56 mJ
Table 6 presents the energy consumption of a CogniSNN with OR Gate, a CogniSNN with ADD operation and a chain-like SNN with OR Gate (which is chain-like connected by the same number of ResNodes), when testing a single sample from DVS-Gesture, CIFAR10-DVS, N-Caltech101 at different time steps.
It is evident that the OR Gate operation consistently consumes less energy than the ADD operation. This is primarily because the OR Gate maintains the output in spike form, thereby preventing infinite amplification and reducing the firing rate of neurons in deeper layers. Additionally, CogniSNN with RGA sometimes consumes slightly more energy than the chain-like SNN, which is intuitive given that the RGA involves a greater computational load. However, this energy consumption gap is acceptable, as the RGA also provides significant performance gains.
@@ 6 Conclusion
This paper validates the feasibility of RGA-based SNNs for modeling the inherent depth scalability and path plasticity in the brain and develops a neuroscience-inspired CogniSNN using the proposed ResNode, the OR gate-driven skip connection mechanism, and the critical path-based LwF algorithm. The results are surprising: Compared to traditional SNNs, the CogniSNN is competitive in image recognition and continual learning. This research opens numerous exciting avenues, including leveraging the scalability of CogniSNN for continual learning by automatically adjusting its structure, or optimizing the graph structure to enhance performance.
7 Acknowledgements
This study is supported by Northeastern University, Shenyang, China (02110022124005) and Guangdong S&T Program (2021B0909060002).